three consecutive integers whose product is 60
Answer by stanbon(75887)

three consecutive integers whose product is 60
-----------------
1st: x-1
2nd: x
3rd: x+1
------------------
Equation:
product = 60
(x-1)x(x+1) = 60
x(x^2-1) = 60
x^3-x-60 = 0
x = 4
---
1st:3
2nd:4
3rd:5
===========
Cheers,
Stan H.
===========